\begin{thebibliography}{99}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\bibitem{And} 
I. Anderson, {\it Combinatorics of finite sets}, corrected reprint of the 1989 edition, 
Dover, Mineola, NY, 2002.

\bibitem{BS07}
J. Bell, M. Skandera,
Multicomplexes and polynomials with real zeros,
Discrete Math. {\bf 307} (2007) 668--682.

\bibitem{Bra} 
P. Br\"and\'en, On operators on polynomials preserving real-rootedness and the Neggers-Stanley conjecture, J. Algebraic Combin. {\bf 20} (2004) 119--130.

\bibitem{BH} 
P. Br\"and\'en and J. Huh, Lorentzian polynomials, Ann. of Math. (2) {\bf 192} (2020) 821--891.

\bibitem{Br89} F. Brenti,
Unimodal Log-Concave and P\'olya Frequency Sequences in Combinatorics,  Mem. Amer. Math. Soc. {\bf 81}, 
Amer. Math. Soc., Providence, 1989.

\bibitem{Cle74} G.~F. Clements, Inequalities concerning numbers of subsets of a finite set, J. Combinatorial Theory Ser. A {\bf 17} (1974), 227--244.

\bibitem{ER94}
P. Edelman, V. Reiner, H-shellings and h-complexes, 
Adv. Math. {\bf 106} (1994) 36--64.

\bibitem{GW96} Garloff, D.G. Wagner, 
Preservation of total nonnegativity under the {H}adamard
              product and related topics,
 In: \textit{Total positivity and its applications ({J}aca, 1994)},
    Math. Appl. {\bf 359},
     pp. {97--102},
 Kluwer Acad. Publ., Dordrecht,1996.

\bibitem{Gash98}
V. Gasharov,
On the Neggers-Stanley conjecture and the Eulerian polynomials,
J. Combin. Theory Ser. A {\bf 82} (1998) 134--146.

\bibitem{HH} 
J. Herzog, T. Hibi, {\it Monomial ideals}, Graduate Texts in Mathematics, {\bf 260}, Springer, London, 2011.

\bibitem{Kat68}
G.O.H. Katona, A theorem of finite sets, In: {\it Theory of Graphs (Proc. Colloq., Tihany, 1966)}, pp. 187--207, Academic Press, New York-London, 1968.

\bibitem{Kru63}
J. Kruskal,
The number of simplices in a complex,
In: Mathematical optimization techniques,
pp. 251-278, University of California Press, Berkeley, CA, 1963.


\bibitem{Maca27}
F. Macaulay,
Some properties of enumeration in the theory of modular systems,
Proc. London Math. Soc. {\bf 26} (1927) 531--555.


 \bibitem{Mur10}
S. Murai,
On face vectors of barycentric subdivisions of manifolds,
SIAM J. Discrete Math. {\bf 24} (2010) 1019--1037.


\bibitem{Slo}
N.J.A. Sloane, The On-Line Encyclopedia of Integer Sequences, \url{http://oeis.org}.

\bibitem{Sta}
  R.P. Stanley, 
  \textit{Combinatorics and commutative algebra},
  Progress in Mathematics, {\bf 41}, Second Edition, 
  Birkh\"auser Boston, Inc., Boston, MA, 1996.

\bibitem{Ste} J.~R. Stembridge, Counterexamples to the poset conjectures of Neggers, Stanley, and Stembridge, Trans. Amer. Math. Soc. {\bf 359} (2007) 1115--1128.
\end{thebibliography}

\documentclass[11pt]{amsart}

\usepackage{amsmath,amsthm,amsfonts,amssymb,latexsym,mathrsfs,color,cite,tikz,cases,url,enumerate}
\usepackage{color}
\usepackage[pagebackref]{hyperref}
\usepackage[capitalize]{cleveref}
\usepackage{enumitem}
\usepackage{verbatim}

\def\lk{\mathrm{lk}}
\def\RR{\mathbb{R}}
\allowdisplaybreaks

  \makeatletter
  \newcommand{\numberlike}[2]{%
     \expandafter\def\csname c@#1\endcsname{%
         \expandafter\csname c@#2\endcsname}%
  }
  \makeatother

  \def\DefaultNumberTheoremWithin{section}

  \theoremstyle{plain}
  \newtheorem{lemma}{Lemma}
     \numberwithin{lemma}{\DefaultNumberTheoremWithin}
     \labelformat{lemma}{Lemma~#1}
  \newtheorem{theorem}{Theorem}
     \numberwithin{theorem}{\DefaultNumberTheoremWithin}
     \numberlike{theorem}{lemma}
     \labelformat{theorem}{Theorem~#1}
  \newtheorem*{maintheorem*}{Corollary of Main Theorem}
  \newtheorem{corollary}{Corollary}
     \numberwithin{corollary}{\DefaultNumberTheoremWithin}
     \numberlike{corollary}{lemma}
     \labelformat{corollary}{Corollary~#1}
  \newtheorem{proposition}{Proposition}
     \numberwithin{proposition}{\DefaultNumberTheoremWithin}
     \numberlike{proposition}{lemma}
     \labelformat{proposition}{Proposition~#1}
  \newtheorem{conjecture}{Conjecture}
     \numberwithin{conjecture}{\DefaultNumberTheoremWithin}
     \numberlike{conjecture}{lemma}
     \labelformat{conjecture}{Conjecture~#1}

  \theoremstyle{definition}
  
     \numberwithin{definition}{\DefaultNumberTheoremWithin}
     \numberlike{definition}{lemma}
     \labelformat{definition}{Definition~#1}
  \newtheorem{question}{Question}
     \numberwithin{question}{\DefaultNumberTheoremWithin}
     \numberlike{question}{lemma}
     \labelformat{question}{Question~#1}
  
     \numberwithin{problem}{\DefaultNumberTheoremWithin}
     \numberlike{problem}{lemma}
     \labelformat{problem}{Problem~#1}
   
  \theoremstyle{remark}
  \newtheorem{remark}{Remark}
     \numberwithin{remark}{\DefaultNumberTheoremWithin}
     \numberlike{remark}{lemma}
     \labelformat{remark}{Remark~#1}
  
     \numberwithin{example}{\DefaultNumberTheoremWithin}
     \numberlike{example}{lemma}
     \labelformat{example}{Example~#1}

     \numberwithin{claim}{\DefaultNumberTheoremWithin}
     \numberlike{claim}{lemma}
     \labelformat{claim}{Claim~#1}
\labelformat{figure}{Figure~#1}
\labelformat{chapter}{Chapter~#1}
\labelformat{appendix}{Appendix~#1}
\labelformat{section}{Section~#1}
\labelformat{subsection}{Subsection~#1}

\title[Real rooted polynomials and $f$-vectors of simplicial complexes]{On a question about real rooted polynomials and $f$-polynomials of simplicial complexes}

\author{Lili Mu} 
\address{School of Mathematics and Statistics, Jiangsu Normal University, Xuzhou 221116, PR China} 
\email{lilimu@jsnu.edu.cn}

\author{Volkmar Welker}
\address{Philipps-Universit\"at Marburg, Fachbereich Mathematik und Informatik, 35032 Marburg, Germany}
\email{welker@mathematik.uni-marburg.de}

\begin{document}

\begin{abstract}
    For a polynomial $f(t) = 1+f_0t+\cdots +f_{d-1}t^d$ with positive integer
    coefficients Bell and Skandera ask if real rootedness of
    $f(t)$ implies that there is a simplicial complex with
    $f$-vector $(1,f_0,\ldots,f_{d-1})$. In this paper we
    discover properties implied by the 
    real rootedness of $f(t)$ in terms 
    of the
    binomial representation $f_i = \binom{x_{i+1}}{i+1}$, $i \geq 0$. We use these to provide a sufficient criterion
    for a positive answer to the question by Bell and Skandera. We also describe two further approaches to the conjecture and use one to verify that some well studied real rooted classical polynomials are f-polynomials.

    Finally, we provide a series of results showing that the set of f-vectors of simplicial complexes is closed under constructions also preserving real rootedness
    of their generating polynomials.
    \end{abstract}
\maketitle
\vspace{-0.5in}

\section{Introduction}

The results of this paper are motivated by a question of Bell and Skandera.

\begin{question}[Bell and Skandera, Question 1.1 in \cite{BS07}] \label{qu:bs} Let $f(t) = 1 + f_0t + \cdots + f_{d-1}t^d$ be a polynomial with positive integer coefficients. If $f(t)$ is real rooted, is $(1,f_0,\ldots, f_{d-1})$ the
$f$-vector $(f_{-1}^ \Delta,\ldots, f_{d-1}^\Delta)$ of a simplicial complex $\Delta$?
\end{question}

Recall that for a simplicial complex $\Delta$ the entries $f_i^\Delta$ of the $f$-vector are the number of $i$-dimensional faces of $\Delta$.
Thus in other words the conclusion of \ref{qu:bs} is that 
$f(t) = f^ \Delta(t)$ is the $f$-polynomial of a simplicial complex $\Delta$. 
We develop three approaches to this question,
which may be of independent
interest, and collect further
evidence for a positive answer to the question. 
The main results of the paper are.
\begin{itemize}
    \item For a real rooted polynomial $f(t) = 1 + f_0t + \cdots + f_{d-1}t^d$ with positive integer coefficients and real
    numbers $x_i \geq i$ such that
    $f(t) = 1 + \displaystyle{\sum_{i=1}^{d}} \binom{x_i}{i}t^i$ we have
    $x_1 \geq \cdots \geq x_d$ (\ref{xk-decreasing}). As a consequence 
    we provide in \ref{cor:hvector} a new proof of \cite[Theorem 4.2]{BS07} showing that
    real rootedness of $f(t)$ implies that the coefficient sequence is the $f$-vector
    of a multicomplex or equivalently the $h$-vector of a Cohen-Macaulay simplicial complex.
    \item If $f(t) = 1 + \displaystyle{\sum_{i=1}^{d}} \binom{x_i}{i} t^i$ is a 
    real rooted polynomial
    with positive integer coefficients, and
    $x_i \geq \lceil x_{i+1} \rceil$ then there is a simplicial complex
    $\Delta$ such that $f(t)$ is the $f$-polynomial of $\Delta$ (\ref{thm:main}).
    \item We exhibit a method based on Murai's
    \cite{Mur10} admissible vectors which can be
    used to demonstrate that classical real rooted
    polynomials (Eulerian, Stirling, derangement)
    are $f$-polynomials of simplicial complexes.
    For Eulerian polynomials this is a known result (see \cite{Gash98, ER94}).
    \item We show in \ref{sec:construction} that classical constructions preserving real rootedness 
    applied to $f$-polynomials $f^\Delta(t)$ of simplicial complexes yield the $f$-polynomial of another 
    simplicial complex.
\end{itemize}

\section{Notation, definitions and basic results}
For a finite set $\Omega$, a simplicial complex over ground set $\Omega$ is
a set of subsets of $\Omega$ such that $\sigma \subseteq \tau \in \Delta$
implies $\sigma \in \Delta$.
We call a $\sigma \in \Delta$ a face of $\Delta$
and the dimension of a face $\sigma$ is $\dim(\sigma) = \# \sigma -1$.
The dimension of $\Delta$ is $\max_{\sigma \in \Delta} \dim(\sigma)$.
If $\Delta$ is a $d$-dimensional simplicial complex then the
vector $f^\Delta = (f_{-1}^ \Delta,\ldots, f_{d}^ \Delta)$ where
$f_i^ \Delta = \# \big\{\, \sigma \in \Delta~\big|~\dim(\sigma) = i\,\big\}$
is called the $f$-vector of $\Delta$
and $f^ \Delta(t) = \displaystyle{\sum_{i=0}^{d+1}} f_{i-1}^ {\Delta} t^ {i}$ is called the $f$-polynomial of $\Delta$. 
Expanding $(1-t)^d \cdot f^ \Delta\big(\,\frac{1}{1-t}\,\big) = h^ \Delta(t) = \displaystyle{\sum_{i=0}^ {d+1} h_i^ \Delta t^ {i}}$ yields the $h$-polynomial
with coefficient sequence $h^\Delta = (h_0^ \Delta,\ldots, h_{d+1}^ \Delta)$ the $h$-vector of $\Delta$.
For a face $F\in \Delta$, denote by $\lk_{\Delta}(F)=\big\{\,G\,:\, F\cup G\in \Delta, F \cap G = \emptyset\,\big\}$ the link of $F$ in $\Delta$.

It is well-known that the $f$-vector of a simplicial complex can be characterized in terms of functions based upon the following expression of a positive integer
$m$ as a sum of binomial coefficients.
Given a positive integer $i$, there exists a unique representation of $m$ in the form
$$m=\binom{a_i}{i}+\binom{a_{i-1}}{i-1}+\cdots+\binom{a_j}{j},$$
where $a_i>a_{i-1}>\cdots>a_j\geq j\geq 1$.
The refer to this representation of $m$ as the $i$\textsuperscript{th} binomial expansion of $m$.
Define the functions $(\mu_i)_{i\geq 1}$ and $(\kappa_i)_{i\geq 1}$  on $\mathbb{N}$ by
$$\mu_i(m)=\binom{a_i}{i-1}+\binom{a_{i-1}}{i-2}+\cdots+\binom{a_j}{j-1}$$
and 
$$\kappa_i(m)=\binom{a_i-1}{i-1}+\binom{a_{i-1}-1}{i-2}+\cdots+\binom{a_j-1}{j-1}.$$
The characterization of $f$- and $h$-vectors of simplicial complexes are due to Kruskal\cite{Kru63}, Katona\cite{Kat68} independently and Macaulay \cite{Maca27} respectively. Note that in combinatorial terms Macaulay's result characterizes $f$-vectors of multicomplexes which is equivalent to being the $h$-vector of a Cohen-Macaulay simplicial complex, see e.g. \cite{Sta} for 
the equivalence and the definition of Cohen-Macaulay simplicial complexes.

\begin{theorem}[Kruskal-Katona, Macaulay]\label{f-h}
Let $f=(1, f_0,\dots,f_{d-1})$ be a vector of positive integers. Then
\begin{itemize}
  \item $f$ is an $f$-vector of a simplicial complex if and only if $\mu_{i+1}(f_i)\leq f_{i-1}$ for $0\leq i\leq d-1$.
  \item $f$ is an $h$-vector of a Cohen-Macaulay simplicial complex if and only if $\kappa_{i+1}(f_i)\leq f_{i-1}$ for $0\leq i\leq d-1$.
\end{itemize}
 
\end{theorem}

\section{Real rooted polynomials with positive integer coefficients}

\subsection{First approach} As usual we consider the binomial coefficient $\binom{n}{k}$ as
a polynomial function in $n$, so that for $x \in \mathbb{R}$ we have
$$\binom{x}{k}=\frac{x(x-1)\cdots (x-k+1)}{k!}.$$
We also write $(x)_k$ for the denominator $x(x-1) \cdots (x-k+1)$.
Note, that the classical recurrence relation $\binom{x}{k}+\binom{x}{k-1}=\binom{x+1}{k}$
still holds and that for $0 < k \in \mathbb{N}$ and any real number 
$y \geq 0$ there is a unique real number $x \geq k-1$ such that $y=\binom{x}{k}$.
Thus for any polynomial $f(t)= 1 + \displaystyle{\sum_{i=1}^d} y_it^i$ with coefficients $y_i > 0$ there
are uniquely defined real numbers $x_i \geq i-1$ such that such that
$f(t)= 1+\sum_{i=1}^d \binom{x_i}{i}t^i$. We call $(x_1,\ldots,x_d)$ the binomial
representation of $f(t)$ or its coefficient sequence. Before we can formulate our first main results we need the following definition. We say that a polynomial
$a_0 + a_1t+\cdots +a_dt^ d$ of degree $d$ is ultra log-concave if
$$\Big[ \,\frac{a_i}{\binom{d}{i}}\,\Big]^ 2
\geq \frac{a_{i-1} a_{i+1}}{\binom{d}{i-1}\binom{d}{i+1}}$$ for $1 \leq i \leq d-1$.

\begin{theorem}\label{xk-decreasing}
Let $f(t) \in \mathbb{R}_{\geq 0}[t]$ with $f(t)=1+\displaystyle{\sum_{i=1}^ d}\binom{x_i}{i}t^{i}$ be an ultra log-concave polynomial and $x_d \geq d$. 
Then  
$x_1\geq x_2\geq\cdots\geq x_{d}$.
\end{theorem}

\begin{proof}
By ultra log-concavity we have 
\begin{align}\label{NI}
\left[\frac{\binom{x_i}{i}}{\binom{d}{i}}\right]^2\geq \frac{\binom{x_{i-1}}{i-1}}{\binom{d}{i-1}}\frac{\binom{x_{i+1}}{i+1}}{\binom{d}{i+1}} \text{ for } 2 \leq i \leq d-1
\end{align}
and \begin{align} \label{N1}
\left[\frac{\binom{x_1}{1}}{\binom{d}{i}}\right]^2 \geq \frac{\binom{x_2}{2}}{\binom{d}{2}}.
\end{align}

From \eqref{NI} and \eqref{N1} it is not hard to derive that
$$\binom{x_i}{i}\geq \left[ \binom{x_d}{d}\right]^{i/d}\binom{d}{i}.$$

Since $x_d \geq d$ we have $\left[ \binom{x_d}{d}\right]^{i/d} \geq 1$
and hence $\binom{x_i}{i}\geq \binom{d}{i}$ and $x_i \geq d$ for all $i$.

The proof of the theorem is now by induction on $i$.
For the case $i=1$, by \eqref{N1}, we have $(d-1)x_1^2\geq dx_2(x_2-1)$.
Assume $x_1< x_2$ then
$$(d-1)x_1^2<(d-1)x_2^2\leq dx_2(x_2-1),$$
which contradicts \eqref{N1} and $x_1\geq x_2$ follows.

For the induction step let $i \geq 2$ and assume that $x_{i}\leq x_{i-1}$.
By \eqref{NI},
$$\frac{(x_i)_i^2}{(d)_i^2}\geq \frac{(x_{i-1})_{i-1}}{(d)_{i-1}}\frac{(x_{i+1})_{i+1}}{(d)_{i+1}}.$$
By the induction hypothesis,
$$\frac{(x_i)_i^2}{(d)_i^2}\geq \frac{(x_{i})_{i-1}}{(d)_{i-1}}\frac{(x_{i+1})_{i+1}}{(d)_{i+1}},$$
which equivalent to
\begin{align}\label{xi}
(d-i)\cdot (x_i)_i\cdot (x_i-i+1)\geq (d-i+1)\cdot (x_{i+1})_{i+1}.
\end{align}
Assume that $x_i<x_{i+1}$, then
$$(d-i)\cdot (x_i)_i\cdot (x_i-i+1)< (d-i)\cdot (x_{i+1})_i\cdot (x_{i+1}-i+1)\leq (d-i+1)\cdot (x_{i+1})_{i+1},$$
where the last inequality follows from $d\leq x_{i+1}$.
This contradicts \eqref{xi} and the proof is complete.
\end{proof}

By \cite[Example 2.3]{BH} and classical results by Newton the assumption of 
\ref{xk-decreasing} is satisfied in the following two cases.

\begin{corollary}
    \begin{itemize}
        \item[(i)] Let $f(t) \in \mathbb{R}_{\geq 0}[t]$ with $f(t)=1+\displaystyle{\sum_{i=1}^ d}\binom{x_i}{i}t^{i}$ and $x_d \geq d$. If $f(t)$ is real rooted then 
$x_1\geq x_2\geq\cdots\geq x_{d}$.
        \item[(ii)] Let $f(t,s) = 
        s^d + \displaystyle{\sum_{i=1}^ d}\binom{x_i}{i}t^{i}s^ {d-i}$ and $x_d \geq d$.
        If $f(s,t)$ is Lorentzian then 
$x_1\geq x_2\geq\cdots\geq x_{d}$.

    \end{itemize}
\end{corollary}

An immediate consequence of \ref{xk-decreasing} is the following result which is the main result in \cite{BS07}, where it is formulated in terms of
$f$-vectors of multicomplexes.

\begin{corollary} \label{cor:hvector}
Let $f(t)=1+\displaystyle{\sum_{i=1}^ d }f_{i-1}t^{i}\in \mathbb{N}(t)$ be a polynomial with only real zeros and
$f_i=\binom{x_{i+1}}{i+1}$. Then $f(t)$ is the $h$-polynomial of a Cohen-Macaulay simplicial complex.
\end{corollary}

\begin{proof}
Suppose
$$f_i=\binom{x_{i+1}}{i+1}=\binom{a_{i+1}}{i+1}+\binom{a_{i}}{i}+\cdots\binom{a_j}{j},$$
then by \ref{xk-decreasing}, we have $a_{i+1}\leq x_{i+1}\leq x_{i}$.
$$\kappa_{i+1}(f_i)=\binom{a_{i+1}-1}{i}+\binom{a_{i}-1}{i-1}+\cdots+\binom{a_j-1}{j-1}\leq \binom{a_{i+1}}{i}\leq \binom{x_{i}}{i}=f_{i-1}.$$
By \ref{f-h}, $f(t)$ is the $h$-polynomial of a simplicial complex.
\end{proof}

Now we are position to formulate and prove our main result supporting 
a positive answer to \ref{qu:bs}.

\begin{theorem} \label{thm:main}
Let $f(t)=1+\displaystyle{\sum_{i=1}^ d \binom{x_i}{i}t^i}\in \mathbb{N}(x)$ be a polynomial with only real roots.
If $x_{i-1}\geq \lceil x_i\rceil$ then $f(t)$ is the $f$-polynomial of a simplicial complex.
\end{theorem}

\begin{proof}
Let
$$f_{i-1}=\binom{x_i}{i}=\binom{a_i}{i}+\binom{a_{i-1}}{i-1}+\cdots+\binom{a_j}{j},$$
with $x_i \geq i$ and $a_i > \cdots a_j \geq j \geq $ be the $i$\textsuperscript{th} binomial expansion of $f_{i-1}$.
If follows that $a_i\leq x_i< a_i+1$.
If $a_i= x_i$, then $f_{i-1}=\binom{x_i}{i}=\binom{a_i}{i}$.
$$\mu_i(f_{i-1})=\binom{x_i}{i-1}=\binom{\lceil x_i\rceil}{i-1}\leq \binom{x_{i-1}}{i-1}=f_{i-2}.$$
If $a_i< x_i<a_i+1$ then $a_i+1=\lceil x_i\rceil \leq x_{i-1}$.
Hence $$\mu_i(f_{i-1})=\binom{a_i}{i-1}+\binom{a_{i-1}}{i-2}+\cdots+\binom{a_j}{j-1}\leq \binom{a_i+1}{i-1}\leq \binom{x_{i-1}}{i-1}=f_{i-2}.$$
By \ref{f-h}, $f(t)$ is the $f$-polynomial of a simplicial complex.
\end{proof}

Data suggests that with this approach the
results from \ref{cor:euler} can be deduced.
Nevertheless, we we not able to verify the assumptions
of \ref{thm:main} for any of the sequences
treated in the corollary.

\begin{remark}
Note that the assumptions of \ref{thm:main} are not always satisfied. 
For example, the polynomial $f(t)=1+4t+5t^2+2t^3 = (x+1)^ 2(2x+1)$
has only real zeros,
but when expressed as $f(t) = 1+\binom{x_1}{1}t + \binom{x_2}{2} +
\binom{x_3}{3}t^ 3$
then $x_1 = 4$, $x_2 = \frac{1}{2} (1+\sqrt{41}) = 3.701562118 \cdots$
and $x_3 = \frac{1}{3}\,\sqrt [3]{162+3\,\sqrt {2913}}+{\frac {1}{\sqrt [3]{162+3\,\sqrt 
{2913}}}}+1 = 3.434841368\cdots$.
It follows that $$x_{2} = 3.701562118 \not\geq 4 = \lceil 3.434841368\cdots \rceil = \lceil x_3 \rceil.$$ 
On the other hand $(1,4,5,2)$ is the $f$-vector of the simplicial complex
given by two triangles glued along an edge.
\end{remark}

\begin{remark}
$f(t)= 1+4t+6t^2+3t^3 = 1+ \binom{4}{1}t + \binom{4}{2}t^2+\binom{3.xxxx}{3}t^3$
is not real rooted. It is the $f$-polynomial of the boundary of the tetrahedron with 
the interior of one triangle removed. On the other hand $x_i \geq \lceil x_{i-1} \rceil$
is still satisfied.
\end{remark}

\begin{remark}
Note that $f$-polynomial with $x_{i-1}\geq \lceil x_i\rceil$ can not get the real rootedness even not the log-concave. For example, the polynomial
$f(t) =1+10t+3t^2+t^3=1+\binom{10}{1}t+\binom{3}{2}t^2+\binom{3}{3}t^3$ is the $f$-polynomial of
a triangle together with $7$ isolated vertices. Clearly, $10 \geq 3\geq 3$ but
$9 = 3^2 \not\geq 1 \cdot 10 = 10$.
\end{remark}


\subsection{Second approach} Next we provide a second approach to a positive answer for
\ref{qu:bs}. 
Let 
$f(t) = 1 + \displaystyle{\sum_{i=1}^ d} f_{i-1} t^ i \in \mathbb{N}[t]$ and
$$f_{i-1} = \binom{x_i}{i} = \binom{a_i}{i}+\binom{a_{i-1}}{i-1}+\cdots+\binom{a_j}{j}, \text{ for } x_i \geq i \text{ and } a_i > \cdots > a_j \geq j \geq 1$$ the $i$\textsuperscript{th} binomial expansion of $f_{i-1}$.
For $i \geq 1$ we define $g_i$ and $h_{i-1}$ by
 $$g_i =\binom{a_i-1}{i}+\binom{a_{i-1}-1}{i-1}+\cdots+\binom{a_j-1}{j}$$
and $$h_{i-1} =\binom{a_i-1}{i-1}+\binom{a_{i-1}-1}{i-2}+\cdots+\binom{a_j-1}{j-1}.$$
It is easily seen that $h_0 = 1$ and $f_{i-1} = g_i + h_{i-1}$. In particular, we have 
$$ f(t) = 1+\sum_{i=1}^ d  f_{i-1}t^ i = \Big(\,1 + \sum_{i=1}^d 
g_i t^ i \,\Big)+ t \Big(\,1 + \sum_{i=2}^ d h_{i-1} t^{i-1}\,\Big).$$
We call the decomposition $f(t) = g(t)+th(t)$ with
$g(t) =1+\sum_{i=1}^d  g_it^ i$ and
$h(t) = 1+\sum_{i=1}^{d-1}  h_it^ i$ the recursive decomposition
of $f(t)$.

We will show that the a positive answer to the following conjecture implies a positive answer to \ref{qu:bs}.


\begin{conjecture} \label{con:second}
Let $f(t)=1+\displaystyle{\sum_{i=1}^ d}  f_{i-1} t^i\in \mathbb{N}[t]$ be a polynomial with only real zeros
and $f(t) = g(t)+th(t)$ its recursive decomposition with 
 $g(t) = 1+\displaystyle{\sum_{i=1}^d} g_it^i$ and $h(t) = 1+\displaystyle{\sum_{i=1}^{d-1}} h_it^i$ then
 $h_i\leq g_i$.
\end{conjecture}

\begin{proposition} If \ref{con:second} holds then 
\ref{qu:bs} has a positive answer.
\end{proposition}
\begin{proof}
Let 
$$f_{i-1} = \binom{a_i}{i}+\binom{a_{i-1}}{i-1}+\cdots+\binom{a_j}{j}, \text{ for }  a_i > \cdots a_j \geq j \geq 1,$$ and 
$$f_{i} = \binom{b_{i+1}}{i+1}+\binom{b_{i}}{i}+\cdots+\binom{b_r}{r}, \text{ for } b_{i+1} > \cdots > b_r \geq r \geq 1$$
the $i$\textsuperscript{th} binomial expansion of $f_{i-1}$ and the $(i+1)$\textsuperscript{st} binomial expansion of $f_i$.
Then  $$g_i=\binom{a_i-1}{i}+\binom{a_{i-1}-1}{i-1}+\cdots+\binom{a_j-1}{j}$$
and $$h_i=\binom{b_{i+1}-1}{i}+\binom{b_{i}-1}{i}+\cdots+\binom{b_r-1}{r-1}.$$
It follows that $h_i\leq g_i$ if and only if 
 $(b_{i+1},b_{i},\dots,b_r)\leq_{lex} (a_i,a_{i-1},\dots,a_j)$ in the lexicographic order $\leq_{lex}$.
 Note that $$\mu_{i+1}(f_{i}) = \binom{b_{i+1}}{i}+\binom{b_{i}}{i-1}+\cdots+\binom{b_r}{r-1}.$$
 If  $(b_{i+1},b_{i},\dots,b_r)\leq_{lex} (a_i,a_{i-1},\dots,a_j)$ then we have $\mu_{i+1}(f_{i})\leq f_{i-1}$.
 The assertion now follows.

\end{proof}

Based on experiments suggesting a positive answer we ask the following question.

\begin{question}\label{que:second}
    Let $f(t)=1+\displaystyle{\sum_{i=1}^d} f_{i-1} t^i\in \mathbb{N}[t]$ be a polynomial with only real zeros
and $f(t) = g(t)+th(t)$ its recursive decomposition then
 $g(t)$ and $h(t)$ both have only real zeros.
\end{question}

\subsection{Third approach} Another technique to prove that a given vector is an $f$-vector of a simplicial complex is the use of admissible vectors, as introduced by Murai in \cite{Mur10}.
Let $f=(1, f_0,\dots,f_{d-1})\in \mathbb{Z}^{d+1}$ be the $f$-vector of a simplicial complex.
We say that a vector $\alpha=(0,1, \alpha_0,\dots,\alpha_{d-2})\in \mathbb{Z}^{d+1}$ is a {\it basic admissible vector} of $f$ if
\begin{itemize}
  \item[(BA1)] $(1, \alpha_0,\dots,\alpha_{d-2}) \in \mathbb{Z}^{d}$ is an $f$-vector of a simplicial complex;
  \item[(BA2)] $f_i\geq \alpha_i$ for $i=0,1,\dots, d-2$.
\end{itemize}
A vector $\beta\in \mathbb{Z}^{d+1}$ is {\it admissible} to $f$ if there exists a sequence of vectors $\beta_1,\dots,\beta_t\in\mathbb{Z}^{d+1}$
such that \begin{itemize}
            \item[(A1)] $\beta=\beta_1+\cdots+\beta_t$;
            \item[(A2)] $\beta_i$ is a  basic admissible vector of $f+\beta_1+\cdots+\beta_{i-1}$ for $i=1,2,\dots,t$.
          \end{itemize}

\begin{lemma}[Lemma 3.1 (1)(3)\cite{Mur10}]\label{admissible}
Let $f=(1, f_0,\dots,f_{d-1})\in \mathbb{Z}^{d+1}$ be the $f$-vector of a simplicial complex
and let $\alpha=(0,1, \alpha_0,\dots,\alpha_{d-2}), \beta=(0,1,\beta_0,\ldots, \beta_{d-2}) \in \mathbb{Z}^{d+1}$ be admissible to $f$. Then 
\begin{itemize}
    \item [(i)] $f+\alpha$ is the $f$-vector of a simplicial complex.
    \item[(ii)] $\alpha+\beta$ is admissible to $f$.
    \end{itemize}
\end{lemma}

The next proposition will be key to the proof that some real rooted classical polynomials are f-polynomials. The result is also an addition
to the list of results in \ref{sec:construction}.

\begin{proposition}\label{thm:f+f'}
Let $f(t)$ be an $f$-polynomial of a simplicial complex. Then $f(t)+tf'(t)$ is an $f$-polynomial of a simplicial complex.
\end{proposition}

\begin{proof}
Suppose that $f(t)=1+\displaystyle{\sum_{i=1}^{d}}f_{i-1}t^i$ is an $f$-polynomial of a simplicial complex $\Delta$.
Then $(1, f_0,\dots,f_{d-1})\in \mathbb{Z}^{d+1}$ is the $f$-vector of $\Delta$.
Let $\beta=(0,f_0, 2f_1,\dots,df_{d-1})\in \mathbb{Z}^{d+1}$.
Then by \ref{admissible}, for showing that $f(t)+tf'(t)$ is an $f$-polynomial of a simplicial complex it suffices to show
 $\beta$ is an admissible vector of $f$.
For each vertex $v_i$, $i=1,\ldots, f_0$, of 
$\Delta$ let
$\beta_i = (0,1,\beta_0^{(i)},\ldots, \beta_{d-2}^{(i)})$, where
$(1,\beta_0^{i},\ldots, \beta_{d-2}^{(i)})$ is
the $f$-vector of the link of $v_i$ in $\Delta$. Then (BA1) is satisfied for all
$\beta_i$ by definition. Since the $(i-1)$-faces
of the link of a vertex are in one to one correspondence to a subset of the $i$-faces
of $\Delta$, condition (BA2) follows for all
$\beta_i$. Indeed it follows for all
$\beta_i$ and for any complex
having an $f$-vector which is coordinatewise 
greater or equal to the one of $\Delta$.
Using this and \ref{admissible}(i),(ii) inductively 
it then follows that 
$f + \beta_1 + \cdots + \beta_{i-1}$ is a basic
admissible vector of $f+\beta_1 + \cdots + 
\beta_i$ for $i=1,\ldots, f_0$. 
In particular, \ref{admissible} implies that
$f+\beta$ is an $f$-vector.

Since a single $i$-face of $\Delta$ gives rise to exactly $i+1$ faces of dimension $i-1$ in the
links of its vertices it follows that
$\beta = (0,f_0,2f_1,\ldots, df_{d-1})$
and the assertion follows.
\end{proof}

The following corollary will turn out to be very useful when analyzing sequences defined through 
recursions.

 \begin{corollary}\label{f-vector}
 Let $(1,f_0,\dots,f_{d-1})$ be the $f$-vector of a simplicial complex. Then $$\mu_{i+1}\big(\,(i+2)f_i\,\big)\leq (i+1)\,f_{i-1} \text{ for } 1\leq i\leq d-1.$$
 \end{corollary}
\begin{proof}
Basic calculus shows that for
$f(t) = 1+\sum_{i=1}^{d}f_{i-1}t^i$ we
have that $f(t)+tf'(t) = 
1+2f_0t+\cdots +(d+1)f_{d-1}$. By  \ref{thm:f+f'} it follows that $(1,2f_0,3f_1,\dots,(i+2)f_i,\dots,(d+1)f_{d-1})$ is the $f$-vector of a simplicial complex.
Then by \ref{f-h}, the result follows.
\end{proof}

We now want to apply these result to show that
Eulerian, Stirling and derangement polynomials are $f$-polynomials. Instead of giving three separate proofs, we consider arrays
$\big[\,T_{d,k}\,\big]_{d,k \in \mathbb{Z}}$ of integers which satisfies the following
conditions:

\begin{itemize}
    \item[(R1)] $T_{1,0}=1$,
    \item[(R2)] $T_{d,k}=0$ for all $d$ and $k$ which do not satisfy $0\leq k\leq d-1$,
    \item[(R3)] For $i=1,2,3$ there are integers $r^{(i)},s^{(i)},t^{(i)}$ such that
    \begin{itemize}
        \item[(i)] $r^{(1)} = 0$ and $t^ {(1)} =1$,
        \item[(ii)] $s^ {(i)} \leq t^{(i)}$, 
        \item[(iii)]  for all $0 \leq k \leq d-1$ and $d \geq 2$ we have $a_{d,k}^{(i)} :=r^{(i)}d+s^{(i)}k+t^{(i)} \geq 0$.
        \end{itemize}
    \item[(R4)] For $0 \leq k \leq d-1$
    and $d \geq 2$ we have
\begin{align}\label{triangle}
T_{d,k}=a_{d,k}^{(1)}T_{d-1,k}+a_{d,k}^{(2)}T_{d-1,k-1}+a_{d,k}^{(3)}T_{d-2,k-1}.
\end{align}
\end{itemize}

For the following choices of parameters we recover
Eulerian, Stirling and derangement numbers
and verify that (R1)-(R4) is satisfied.

\begin{itemize}
  \item {\sf (Eulerian)}  $[T_{d,k}]_{d\geq 1, k\geq 0}$ is the Eulerian triangle,
  with
  $T_{d,k}$ the number of permutations in $S_d$ with $k$ descents (see \cite[A008292]{Slo}) for $a_{d,k}^{(1)}=k+1
  = 0 \cdot d+1\cdot k + 1$, $a_{d,k}^{(2)}=
  d-k = 1 \cdot d + (-1) \cdot k + 0$ and $a_{d,k}^{(3)}=0 = 0 \cdot d + 0 \cdot k + 0$.
  \item {\sf (Stirling)} $[T_{d,k}]_{d\geq 1, k\geq 0}$ is the Stirling triangle of the second kind,
  with $T(d,k)$ the number of set partitions of
  $\{1,\ldots, d\}$ with $k+1$ blocks (see \cite[A008277]{Slo}) for $a_{d,k}^{(1)}=k+1 = 0 \cdot d+1\cdot k + 1$, $a_{d,k}^{(2)}=1= 0 \cdot d + 0 \cdot k + 1$ and $a_{d,k}^{(3)}=0 =0 \cdot d + 0 \cdot k + 0$.
  \item {\sf (Derangements)} $T_{d,k}$ is the number of derangements of $[d]$ having $k$ exceedances (see \cite[A046739]{Slo}) for $a_{d,k}^{(1)}=k+1= 0 \cdot d+1\cdot k + 1$, $a_{d,k}^{(2)}=d-k= 1 \cdot d + (-1) \cdot k + 0$ and $a_{d,k}^{(3)}=1 \cdot d+0 \cdot k +0$.
\end{itemize}

\begin{theorem} \label{thm:T-poly}
Let $[T_{d,k}]_{d,k \in \mathbb{Z}}$ be a triangular array defined satisfying \eqref{triangle}, (R1), (R2), (R3) and (R4).
Then the $d$\textsuperscript{th} row polynomial $T_d(t)=\displaystyle{\sum_{k\geq 0}}T_{d,k}t^k$ is an $f$-polynomial of a simplicial complex.
\end{theorem}

\begin{proof}
Note that by (R2) and (R4) $T_{d,0}= a_{d,0}^{(1)} T_{d-1,0}$ for $d \geq 2$. Now (R1) and (R3)(i) imply $T_{d,0} = 1$ for all $d\geq 1$.

For $d=1$ condition (R1) and (R2) imply that 
$T_1(t) = 1$ is the $f$-polynomial of the simplicial complex $\{ \emptyset \}$. 

For $d=2$ we have by (R1), (R2) and (R3)(iii) that $T_{2,1} = a_{2,1}^{(2)} \geq 0$. And $(1,a_{2,1}^ {(2)})$ is the $f$-polynomial
of the simplicial complex $a_{2,1}^ {(2)}$
simplices of dimension $0$ and the empty set. 

Suppose that $d \geq 3$ and for $n \leq d-1$ 
all $T_n(t)$ are $f$-polynomials 
of simplicial complexes.
By \ref{f-h} we need to show that for $1 \leq k \leq d-1$
we have $\mu_k(T_{d,k)} \leq T_{d,k-1}$.

We will repeatedly use the fact 
that $\mu_k$ is a subadditive function on
non-negative integers. For that we first extend
$\mu_k$ to all
non-negative integers by setting $\mu_k(0) = 0$. It is easily checked that the  subadditivity of $\mu_k$ on positive integers (see \cite[Theorem 7.4.2]{And} or \cite{Cle74}) extends. We mark arguments using the subadditivity by $(sub)$.

\medskip

\noindent {\sf Claim:} 
For $d \geq 3$ and $0 \leq k \leq d-1$ we have 
\begin{itemize}
    \item[(1)] $\mu_k\left(a_{d,k}^{(1)}T_{d-1,k}\right)
\leq a_{d,k-1}^{(1)}T_{d-1,k-1}$,
    \item[(2)] $\mu_k\left(a_{d,k}^{(2)}T_{d-1,k-1}\right)
\leq a_{d,k-1}^{(2)}T_{d-1,k-2}$,
 \item[(3)] $\mu_k\left(a_{d,k}^{(3)}T_{d-2,k-1}\right)
\leq a_{d,k-1}^{(3)}T_{d-2,k-2}$.
\end{itemize}

\medskip

We prove the inequality (1) of the claim. The others then follow by similar reasoning.

\noindent {\sf Case:} $s^{(1)}\leq 0$.

By induction on $d-1$ and since by
 (R3)(iii) $a_{d,k}^{(1)}\geq 0$, we have
 \begin{align*}
\mu_k\left(a_{d,k}^{(1)}T_{d-1,k}\right)&=
\mu_k\left((r^{(1)}d+s^{(1)}k+t^{(1)})\,T_{d-1,k}\right)\\
&\overset{(sub)}{\leq} (r^{(1)}d+s^{(1)}k+t^{(1)})\,\mu_k(T_{d-1,k})\\
&\leq (r^{(1)}d+s^{(1)}(k-1)+t^{(1)})\,T_{d-1,k-1}\\
&=a_{d,k-1}^{(1)}T_{d-1,k-1}.
\end{align*}

\noindent {\sf Case:} $s^{(1)}>0$.

By $s^{(1)}\leq t^{(1)}$ and $a_{d,k}^{(1)}\geq 0$,
\begin{align*}
\mu_k\left(a_{d,k}^{(1)}\,T_{d-1,k}\right)&=
\mu_k\left((r^{(1)}d+s^{(1)}k+t^{(1)})\,T_{d-1,k}\right)\\
&=\mu_k \left((r^{(1)}d+s^{(1)}(k+1)+t^{(1)}-s^{(1)})\,T_{d-1,k}\right)\\
&\overset{(sub)}{\leq} \mu_k \left(s^{(1)}(k+1)\,T_{d-1,k}\right)+ \mu_k \left((r^{(1)}d+t^{(1)}-s^{(1)})\,T_{d-1,k}\right) \\
&{\leq} (r^{(1)}d+s^{(1)}k+t^{(1)}-s^{(1)})\,T_{d-1,k-1}\\
&=a_{d,k-1}^{(1)}T_{d-1,k-1},
\end{align*}
where the last inequality follows by induction on $d-1$ and
\ref{f-vector}. This completes the proof of the claim. 
Using (R4) and (R3)(iii) we conclude
\begin{align*}
\mu_k(T_{d,k})&=\mu_k\left(a_{d,k}^{(1)}T_{d-1,k}+a_{d,k}^{(2)}T_{d-1,k-1}+a_{d,k}^{(3)}T_{d-2,k-1}\right)\\
&\overset{(sub)}{\leq} \mu_k\left(a_{d,k}^{(1)}T_{d-1,k}\right)+\mu_k\left(a_{d,k}^{(2)}T_{d-1,k-1}\right)+\mu_k\left(a_{d,k}^{(3)}T_{d-2,k-1}\right) \\
&\overset{\text{(Claim)}}{\leq} a_{d,k-1}^{(1)}T_{d-1,k-1}+a_{d,k-1}^{(2)}T_{d-1,k-2}+a_{d,k-1}^{(3)}T_{d-2,k-2}\\
&= T_{d,k-1}.
\end{align*}

The assertion now
follows.
\end{proof}

Since we have already verified conditions
(R1)-(R4) for these classical polynomials 
the following corollary is an immediate consequence of \ref{thm:T-poly}.
The case of Eulerian polynomial is
known by work of Edelman and Reiner \cite{ER94} and Gasharov \cite{Gash98}. For the other two we are not aware of any reference.

\begin{corollary} \label{cor:euler}
The Eulerian polynomial, the Stirling polynomial of the second kind and the derangement polynomial are $f$-polynomials of simplicial complexes.
\end{corollary}

The conjecture by Neggers and Stanley states
that the $W$-polynomial $W_P(t)$, also called $P$-Eulerian 
polynomial, of a poset $P$ is real rooted
(see for example \cite{Br89} for definitions
and the conjecture). 
If $P$ is the $n$-element antichain then
its $P$-Eulerian polynomial is the 
$n$\textsuperscript{th} Eulerian polynomial.
$P$-Eulerian polynomials have constant coefficient $1$ if and only if the poset if naturally labeled. This is the case for which
Neggers formulated the conjecture originally. The general Neggers-Stanley conjecture was 
shown to be false by Br\"and\'en \cite{Bra}
and in the naturally labeled case by 
Stembridge \cite{Ste}. 
Assuming the answer to \ref{qu:bs} is positive, then for a naturally labeled poset
whose $P$-Eulerian polynomial $W_P(t)$ is real rooted there is a simplicial complex whose $f$-polynomial is $W_P(t)$. 
Checking the
explicit counterexamples from \cite{Ste} 
conjecture one finds that all satisfy the
Kruskal-Katona conditions from  \ref{f-h}.
This makes us raise the following question.

\begin{question} Let $P$ be a naturally
labeled poset. Is there always a simplicial
complex $\Delta_P$ for which the $f$-polynomial is the $W$-polynomial $W_P(t)$ of $P$?
\end{question}

The result by Edelman and Reiner \cite{ER94} and Gasharov \cite{Gash98}, reproved in 
\ref{cor:euler}, answers the question positively in case $P$ is an antichain. 

Since $f$-vectors are not always unimodal,
the question is independent with respect to
the still open questions about unimodality 
and log-concavity of $W_P(t)$.

\section{Preserving real rootedness preserves the \texorpdfstring{$f-$}~vector property} \label{sec:construction}

In this section we collect results which show that well studied
constructions on polynomials, that preserve real rootedness, also preserve
the property of being the $f$-polynomial of a simplicial 
complex. 

We start with the dilation of the variable. Elementary analysis shows that if $f(t)$ is a polynomial
with only real roots then same holds for $f(ct)$ for
any $c \in \RR$.

\begin{proposition}
Let $f(t)$ be an $f$-polynomial of a simplicial complex. Then for any positive $c \in \mathbb{Z}$ the polynomial $f(ct)$ is an $f$-polynomial of a simplicial complex.
\end{proposition}

\begin{proof}
Let $\Delta$ be a simplicial complex on 
vertex set $[d]$. We construct a simplicial
complex $\Delta^c$ on vertex set $[d]\times [c]$.
Define $\pi$: $[d]\times [c]\rightarrow [d]$ by $\pi\big(\,(i,j)\,\big)=i$.
For $\sigma \subseteq [d]\times [c]$ we set 
$\sigma \in \Delta^c$ if and only if 
$\pi(\sigma) \in \Delta$. Thus for any $i$-simplex $\sigma\in \Delta$ there are $c^{i+1}$
simplices in $\Delta^ c$.  
It follows that if $f(x)$ is the $f$-polynomial of $\Delta$ then $f(cx)$ is the $f$-polynomial of $\Delta^c$.
\end{proof}

It is easily seen that if $(a_n)_{n \geq 0}$ is
a P\'olya-frequency sequence, then so 
is $(a_{kn})_{n \geq 0}$ for any natural number $k \geq 1$
(see e.g. \cite{Br89}). In particular by the Aissen-Edrei-Schoenberg-Whitney Theorem, if 
$f(t) = 1+f_0t+ \cdots +f_{d-1}t^{d}$ is real rooted
then so is $1+f_kt+f_{2k}t^ 2+\cdots$. 
For proving the following result we will resort
to methods from combinatorial commutative algebra 
and the theory of Gr\"obner bases. In the proof
we will use these methods freely and 
refer the reader to \cite{HH} for further
details.

\begin{proposition}
Let $f=(1, f_0,\ldots,f_{d-1})$ be the $f$-vector of a simplicial complex.
Then for any $k \geq 1$ the vector $(1, f_k,f_{2k},\ldots)$ is the $f$-vector of a simplicial complex.
\end{proposition}
\begin{proof}
For this result we use the fact that $(1,f_0,\ldots, f_{d-1})$ is the $f$-vector of a simplicial complex if and only
if for any field $k$ there is a number $n$ and a squarefree monomial 
ideal $I$ in $S = k[x_1,\ldots, x_n]$ such that
the Hilbert series of $R= S/\big(\,I+(x_1^ 2,\ldots, x_n^ 2)\,\big)$ 
is $1+f_0t +\cdots +f_{d-1}t^ {d}$.  
In the latter case we have a direct sum decomposition
as $k$-vectorspaces
$R = A_0\oplus A_1\oplus \cdots \oplus A_{d-1}$ where $A_i$ is the subspace generated by the 
cosets of the homogeneous polynomials of degree $i$
and is of $k$-dimension $f_i$. 
Now $R^{\langle k \rangle} = A_0 \oplus A_k
\oplus A_{2k} \oplus \cdots$ is known as the
$k$\textsuperscript{th} Veronese algebra of $A$. 
We can write $R^{\langle k \rangle}$ as the 
image
of $$T = k\big[\,y_{\ell_1,\ldots,\ell_n}~\big|~\ell_1+\cdots +
\ell_n = k \text{ and } \ell_i \geq 1\, \big]$$
by the homomorphism $\phi$ which sends $y_{\ell_1,\ldots,\ell_n}$ to $x_1^{\ell_1}\cdots x_n^{\ell_n}$. 
Since $x_i^2 = 0$ in $R$ it follows that
all $y_{\ell_1,\ldots,\ell_n}^2$ lie in the kernel of 
$\phi$. Let $M$ be the ideal generated by
the $y_{\ell_1,\ldots,\ell_n}^2$ and
$J$ be such that $J+M = \ker(\phi)$. 
Let $\preceq$ be a term order. Then 
$T/(M+J)$ and $T/{\mathrm in}_\preceq{(M+J)}$ share
the same Hilbert series, where ${\mathrm in}_\preceq(M+J)$
denotes the initial ideal of $M+J$.
Since $M$ is a monomial ideal we clearly have
$M \subseteq {\mathrm in}_\preceq(M+J)$. Since ${\mathrm in}_\preceq(M+J)$ also is a monomial ideal it follows
that ${\mathrm in}_\preceq(J+M) = J' + M$ for a 
squarefree monomial ideal $J'$. 
From this we deduce that 
$(f_{-1},f_k,f_{2k},\ldots)$ is an $f$-vector.
\end{proof}

The fact that products of real rooted polynomials are real rooted is trivial. Also
the first part of the following proposition, which is the $f$-vector analog of this statement is well known.
The real rootedness of the Hadamard product
of two real rooted polynomials is again a known face (see e.g. \cite{GW96}). 

\begin{proposition}
Let $f(t)=1+\displaystyle{\sum_{i=1}^ {d}}  f_{i-1}t^i$ and $g(t)=1+\displaystyle{\sum_{i=1}^{d'}}  g_{i-1} t^i$ be two 
$f$-polynomials of some simplicial complexes. Then the product $f(t)g(t)$ and the Hadamard product $f(t)\circ g(t)=1+\displaystyle{\sum_{i=1}^{\min{d,d'}}} f_{i-1}g_{i-1}t^i$ are $f$-polynomials of some simplicial complexes.
\end{proposition}
\begin{proof}
    Let $\Delta$ be a simplicial complex over
    ground set $\Omega$ with
    $f$-vector $(1,f_0,\ldots,f_{d-1})$ and let 
    $\Delta'$ be a simplicial complex over
    ground set $\Omega'$ with 
    $f$-vector $(1,g_0,\ldots,g_{d'-1})$. 
 We can assume that $\Omega$ and $\Omega'$ are disjoint.
 
    \noindent {\sf Ordinary Product:} 
   The join $\Delta * \Delta'$ of
    $\Delta$ and $\Delta'$ is defined as
    the simplicial complex whose $i$-dimensional
    simplices are unions $\sigma \cup \tau$ for 
    a $j$-dimensional simplex $\sigma \in \Delta$
    and $(i-j-1)$-dimensional simplex $\tau$ of $\Delta'$
    for some $-1 \leq j \leq i$.
    By disjointness of $\Omega$ and $\Omega'$ it follows
    that the number of $i$-dimensional faces of 
    $\Delta*\Delta'$ is $\sum_{j=-1}^i f_jg_{i-j-1}$.
    The latter expression is exactly the coefficient of
    $t^i$ in $f(t)g(t)$. In particular, $f(t)g(t)$ is the $f$-polynomial of a simplicial complex.
    
    \noindent {\sf Hadamard Product:} We assume in addition that $\Omega$ and
    $\Omega'$ are linearly ordered by $<$ and $<'$ respectively. 
    Let $\Delta \cdot \Delta'$ be
    the simplicial complex over
    ground set $\Omega \times \Omega'$ where
    $\{(\omega_1,\omega_1'),\ldots, 
    (\omega_i,\omega_i')\}$ is a simplex in 
    if any only if 
    \begin{itemize}
    \item $\{\omega_1,\ldots,\omega_i\} \in \Delta$
    and $\{\omega_1',\ldots, \omega_i'\}\in \Delta'$
    and
    \item $\omega_1 < \cdots < \omega_i$ and
    $\omega_1' <' \cdots <' \omega_i'$.
    \end{itemize}
    One easily checks that $\Delta \cdot \Delta'$
    is a simplicial complex with $f$-vector
    $(f_{-1}g_{-1},f_0g_0,\cdots)$.
\end{proof}

\section*{Acknowledgments}
We thank Petter Br\"and\'en for informing us about the connection of an 
earlier version of \ref{xk-decreasing} and Lorentzian polynomials. 

The first author was partially supported by the National Natural Science Foundation of China (Grant Nos. 12271222), Scientific Research Foundation of Jiangsu Normal University (Grant Nos. 21XFRS019) and the China Scholarship Council.

\end{document}